OpenStudy (anonymous):

FINAL PROBLEM ! Woot woot, I can study for my other subjects AND SLEEP ! Make "q" the subject of A = B/p-q

OpenStudy (anonymous):

make this the subject of a long novel

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

\[A=\frac{B}{p-q}\] \[\frac{1}{A}=\frac{p-q}{B}\] \[\frac{B}{A}=p-q\] \[\frac{B}{A}-p=-q\] \[q=p-\frac{B}{A}\]

OpenStudy (anonymous):

have a good night.

OpenStudy (anonymous):

Thank u so much Sat! N I sure will. :)